Please help!! Been out of school 14 yrs...
Directions: Solve for the indicated variable.
1. 5x + y = 8;y <--- wanting to say do what??

And this one.
Directions:
Tell which equation you would use to isolate a variable. Explain your reasoning.
1. 3x - y = 5
2x + y = 0
I just don't get it!

Answer by stanbon(48569)

Solve for the indicated variable.
1. 5x + y = 8;y <--- wanting to say do what??
Equation:
5x+y = 8
Subtract 5x from both sides to get:
y = -5x+8
-----------------------
And this one.
Directions:
Tell which equation you would use to isolate a variable. Explain your reasoning.
3x - y = 5
2x + y = 0
----------------
Solve the 2nd equation for "y" to get:
y = -2x
The reason for using the 2nd equation is nothing
more than that the result is simpler.
There is no magic involved.
===================
Cheers,
Stan H.